OpenStudy (anonymous):

solve for x z=17+4(x+y)

OpenStudy (wolf1728):

Since there are 3 unknowns, we need 3 equations to solve for all unknowns. We need 2 more equations.

OpenStudy (anonymous):

If your quedtion is value of x in terms of y & z means then the answer is given z=17+4(x+y) z-17=4(x+y) x+y=(z-17)/4 x={(z-17)/4}-y